### v1

# AgentNoiseBench: Benchmarking Robustness of Tool-Using LLM Agents Under Noisy Condition

Abstract
Recent advances in large language models have enabled LLM-based agents to achieve strong performance on a variety of benchmarks. However, their performance in real-world deployments often that observed on benchmark settings, especially in complex and imperfect environments. This discrepancy largely arises because prevailing training and evaluation paradigms are typically built on idealized assumptions, overlooking the inherent stochasticity and noise present in real-world interactions. To bridge this gap, we introduce AgentNoiseBench, a framework for systematically evaluating the robustness of agentic models under noisy environments. We first conduct an in-depth analysis of biases and uncertainties in real-world scenarios and categorize environmental noise into two primary types: user-noise and tool-noise. Building on this analysis, we develop an automated pipeline that injects controllable noise into existing agent-centric benchmarks while preserving task solvability. Leveraging this pipeline, we perform extensive evaluations across a wide range of models with diverse architectures and parameter scales. Our results reveal consistent performance variations under different noise conditions, highlighting the sensitivity of current agentic models to realistic environmental perturbations. Code is available at https://github.com/keven-cyber/agentnoisebench
1 Introduction
Recent advances in large language models (LLMs) (OpenAI, 2025c; Google, 2025c; Team et al., 2025b) have substantially improved their reasoning and tool-use capabilities (Team et al., 2025a; Zeng et al., 2025a; Team et al., 2025c), enabling the development of increasingly capable LLM-based agents. As a result, agentic models have demonstrated strong performance across a wide range of benchmarks (Yao et al., 2024; Barres et al., 2025; He et al., 2025). However, recent studies indicate that many agents experience notable performance degradation when encountering out-of-distribution settings or operating in complex real-world environments (Deng et al., 2023; Zhou et al., 2023; Xue et al., 2025).
We argue that a fundamental gap exists between the assumptions underlying current agent evaluations and the conditions encountered in realistic environments. In particular, most existing benchmarks evaluate agents under idealized assumptions, where instructions are carefully curated, and interactions with the environment are stable and well-controlled (Zeng et al., 2024; Qi et al., 2024; Jin et al., 2025). In contrast, real-world environments are inherently stochastic and imperfect. User interactions exhibit substantial diversity and unpredictability (Gallois et al., 2005; Trippas et al., 2024; Wang et al., 2024), while external tools frequently return noisy, incomplete, or failed outputs due to uncontrollable factors (Vuddanti et al., 2025; Xiong et al., 2025). Despite the prevalence of such noise in practice, existing benchmarks often fail to systematically expose agents to these factors, making robustness under realistic conditions difficult to assess.
Systematically evaluating agent robustness under noise remains non-trivial. Concretely, there are several main challenges: (1) Absence of a systematic taxonomy: agents are exposed to diverse and heterogeneous noise sources in realistic environments, yet existing benchmarks provide no unified framework for characterizing them. (2) Complexity of noise injection: injecting realistic noise in a principled manner is non-trivial, as it requires balancing increased environmental complexity with the preservation of task solvability. (3) Limitations of evaluation protocols: evaluating agent behavior under noise is inherently multi-faceted, while current benchmarks lack comprehensive evaluation protocols to capture such behaviors.
To this end, we introduce AgentNoiseBench, a systematic framework designed to evaluate the noise robustness of agentic models. Specifically, we conduct a systematic analysis of biases observed in real-world environments and identify two primary sources of noise: (1) user-noise, which captures the inherent ambiguity and variability in user interaction patterns, and (2) tool-noise, which simulates execution failures, inconsistent responses, and partial results from external tools. Building on this analysis, we develop an automated pipeline to inject both noise types into existing agentic benchmarks, while enforcing strict constraints to preserve task solvability. Furthermore, we introduce a trajectory-aware evaluation protocol that enables multi-dimensional analysis of agent behavior under noisy conditions.
We instantiate AgentNoiseBench across two representative agentic capabilities—tool use and search—by injecting controlled user-noise and tool-noise into -Bench and VitaBench (Barres et al., 2025; He et al., 2025), as well as HotPot QA (Yang et al., 2018). We then evaluate a broad set of open-source and proprietary models, including both reasoning-oriented models and standard foundation models, spanning diverse architectures and parameter scales. As shown in Figure 1, all evaluated models exhibit consistent performance degradation under both noise sources, with varying degrees of sensitivity. Further analysis indicates that general reasoning ability and environmental robustness are not strongly correlated. By examining the step-wise entropy of agent trajectories, we find that different noise sources affect the entropy minimization process through distinct mechanisms. Finally, systematic failure pattern analysis shows that models across all families are more sensitive to tool-noise than to user-noise, potentially due to the inherent variability and limited controllability of the user simulator.
Overall, the main contributions of this work are:
We conduct a systematic analysis of biases in realistic environments and categorize them into two primary noise sources: user-noise and tool-noise.
We develop an automated pipeline that injects controllable noise into existing agentic benchmarks while preserving task solvability.
We introduce a trajectory-aware evaluation protocol and perform extensive evaluations across a wide range of open-source and proprietary models, yielding key insights into agent robustness.
2 AgentNoiseBench: A Systematic Framework for Robustness Evaluation
In this section, we formally present AgentNoiseBench, a systematic evaluation framework designed to quantify the robustness of LLM-based agents against both random perturbations (Song et al., 2024; Li et al., 2024) and adversarial imperfections (Zhan et al., 2024; Zhang et al., 2025) inherent in real-world deployment.
2.1 Design Principles
Real-world environments are inherently imperfect (Wang et al., 2025a; Siska et al., 2024; Lunardi et al., 2025). To address this evaluation gap, AgentNoiseBench bridges the gap between idealized benchmark evaluations and the complexities of realistic environments through three governing principles:
Realistic Noise Modeling. Rather than injecting arbitrary noise, we model valid noise grounded in real-world interaction logs. Specifically, we categorize noise into user-noise, reflecting user cognitive biases and linguistic variability (Zhang et al., 2024a; Gan et al., 2024; Rivkin et al., 2023), and tool-noise, simulating instability and unreliability in external environments (Xu et al., 2024a; Kokane et al., ; Chen et al., 2025; Zhang et al., 2024b).
Solvability-Preserving Noise Injection. A core premise of our framework is that injected noise should increase task difficulty without rendering the task unsolvable. Accordingly, all perturbations are strictly constrained to preserve the necessary conditions for a valid solution, ensuring that failures can be attributed to agent fragility rather than task invalidity.
Trajectory Consistency Enforcement. Outcome-based metrics are insufficient as they fail to distinguish between robust reasoning and stochastic luck. Therefore, a rigorous evaluation must demand procedural integrity: the agent’s intermediate reasoning steps should remain logically sound and strictly adhere to task requirements, impervious to noise-induced deviations. This ensures that success reflects genuine stability in the decision-making process rather than accidental correctness.
2.2 Benchmark Construction
To ensure a fair and standardized comparison across different agents, our benchmark construction adheres to a unified testbed principle. We employ a data-driven analysis of user–agent interaction logs to derive realistic noise patterns. Furthermore, to guarantee that adversarial challenges remain consistent across all evaluated models, we adopt a proxy-guided generation strategy, ensuring that the difficulty level is controlled by a fixed, standardized noise generation protocol rather than adapting to specific victim models.
2.2.1 Empirical Taxonomy of Noise
We derive a noise taxonomy through systematic analysis of large-scale user–agent interaction logs. Specifically, we combine manual annotation with frequency-based statistics to identify noise patterns that are both prevalent in practice and associated with high interaction costs.
User Noise ():
User noise captures imperfections and variability in human-provided instructions during agent interactions. Specifically, we consider the following categories of user-side perturbations:
Ambiguous Instructions: Instructions that are underspecified or lack explicit constraints.
Inconsistent Instructions: Contradictory or conflicting directives issued across multi-turn interactions.
Redundant Information: The inclusion of excessive irrelevant or low-utility contextual information.
Topic Drift: Shifts in the user’s primary intent or conversational focus across dialogue turns.
Boundary Probing: Queries that deliberately test the limits of the agent’s knowledge or reasoning capabilities.
Tool Noise ():
Tool noise captures the stochasticity and unreliability of external tools, APIs, and environments during task execution. Specifically, we consider the following categories of tool-side perturbations:
Execution Failures: Tool invocations fail to return valid responses due to service-level issues, including unavailable or uninitialized tools, inaccessible external services, or server-side exceptions.
Incomplete Responses: Tools return partial or malformed outputs, such as truncated JSON objects or missing fields required for downstream steps.
Erroneous Outputs: Tools produce responses containing incorrect or internally inconsistent factual information.
Misleading Signals: Tool outputs include misleading or unreliable content that can divert agents toward incorrect reasoning trajectories.
Redundant Information: Tools return excessive irrelevant or low-utility content, unnecessarily increasing context length and processing overhead.
2.2.2 Constrained Adversarial Noise Injection
To generate noise that is strictly valid yet sufficiently challenging, while ensuring fairness across diverse models, we introduce a constrained adversarial injection mechanism. Specifically, we employ a state-of-the-art model (e.g., o3) as a fixed reference agent to derive a universal noise generation system prompt.
We instantiate a strong language model as the noise generator , with its parameters frozen. Optimization is performed exclusively over the parameterized adversarial system prompt . Given an original task input , the generator produces a perturbed task instance conditioned on . Our objective is to identify an optimal prompt strategy that maximally degrades the performance of the reference agent , subject to the strict preservation of task solvability:
Here, quantifies the performance degradation of the reference agent under injected noise, while enforces a solvability constraint to exclude ill-posed task instances. We iteratively update to steer the model toward producing adversarial perturbations that induce maximal failure without violating task validity.
By anchoring the optimization process to a state-of-the-art proxy, we leverage the transferability of adversarial examples: noise capable of disrupting the most robust agent serves as a rigorous, high-bar challenge for all other models. Once identified, is frozen and applied uniformly to generate noise for all evaluated agents. This ensures that the benchmark difficulty remains static and unbiased, preventing adaptive overfitting to specific victim models, thereby systematically constructing the hardest valid evaluation instances.
2.3 Trajectory-Aware Evaluation Protocol
We retain the standard evaluation metrics from VitaBench (He et al., 2025) and -Bench (Barres et al., 2025) to assess final outcome correctness. To rigorously capture procedural integrity, we additionally introduce a Trajectory-Aware Evaluation Protocol. This mechanism scrutinizes the entire interaction history, penalizing agents that arrive at correct answers via corrupted or unsafe reasoning paths.
Trajectory Deviation Detection. Let denote the agent’s reasoning trajectory. Given a task specification and a noise injection plan , we first define a step-wise validity indicator . For each state , a model-based evaluator verifies whether the agent’s behavior remains consistent with the task requirements:
A trajectory is deemed valid only if all intermediate steps maintain consistency. Consequently, the trajectory-level validity is defined as the conjunction of step-wise indicators:
Stability-Gated Success Metric. To strictly enforce robustness, we propose a stability-gated accuracy (SGA) metric. This metric regards a task as successfully solved if and only if the agent produces the correct outcome and maintains a flawless reasoning trajectory:
where denotes the binary success of the final outcome. Under this rigorous standard, any intermediate deviation () collapses the entire metric to zero, effectively filtering out false positive success cases driven by stochastic luck.
| Models | -bench | Vitabench | Search | |||||
| Origin | Tool_Noise | User_Noise | Origin | Tool_Noise | User_Noise | Origin | Tool_Noise | |
| Non-thinking Models | ||||||||
| DeepSeek-V3-0324 | 0.470.02 | 0.310.01 | 0.370.03 | 0.250.02 | 0.170.01 | 0.200.02 | 0.210.01 | 0.220.03 |
| Gemini-2.5-Flash (think off) | 0.560.04 | 0.310.02 | 0.500.01 | 0.260.03 | 0.160.02 | 0.220.01 | 0.280.02 | 0.200.02 |
| Doubao-Seed-1.6 | 0.510.03 | 0.310.02 | 0.450.04 | 0.370.01 | 0.280.03 | 0.320.02 | 0.260.03 | 0.200.02 |
| GPT-4.1 | 0.560.02 | 0.300.01 | 0.500.03 | 0.410.04 | 0.310.02 | 0.330.01 | 0.290.03 | 0.270.04 |
| DeepSeek-V3.1 (w/o thinking) | 0.550.01 | 0.350.03 | 0.450.02 | 0.360.03 | 0.310.01 | 0.300.02 | 0.430.04 | 0.370.01 |
| DeepSeek-V3.2-Exp (w/o thinking) | 0.520.03 | 0.340.02 | 0.490.01 | 0.420.04 | 0.320.03 | 0.320.01 | 0.450.02 | 0.360.04 |
| gemini-2.5-pro (w/o thinking) | 0.590.02 | 0.370.04 | 0.550.03 | 0.490.01 | 0.340.02 | 0.390.04 | 0.360.03 | 0.290.03 |
| gpt-3.5-turbo (w/o thinking) | 0.530.04 | 0.280.01 | 0.460.02 | 0.360.03 | 0.230.01 | 0.260.03 | 0.170.02 | 0.130.01 |
| Qwen3-Max | 0.570.01 | 0.340.03 | 0.510.04 | 0.360.02 | 0.230.03 | 0.230.02 | 0.310.03 | 0.240.02 |
| GLM-4.5 (w/o thinking) | 0.620.03 | 0.420.02 | 0.550.01 | 0.410.04 | 0.290.01 | 0.310.03 | 0.440.02 | 0.360.02 |
| GLM-4.6 (w/o thinking) | 0.740.02 | 0.480.04 | 0.640.03 | 0.400.01 | 0.320.02 | 0.330.04 | 0.470.01 | 0.390.02 |
| LongCat-Flash-Chat | 0.740.04 | 0.480.01 | 0.620.02 | 0.430.03 | 0.340.04 | 0.360.01 | 0.380.03 | 0.300.02 |
| Claude-4-Sonnet (w/o thinking) | 0.650.01 | 0.410.03 | 0.550.04 | 0.440.02 | 0.340.01 | 0.330.03 | 0.380.04 | 0.300.01 |
| Claude-4.5-Sonnet (w/o thinking) | 0.710.03 | 0.480.02 | 0.640.01 | 0.480.04 | 0.400.03 | 0.390.02 | 0.340.01 | 0.290.03 |
| Thinking Models | ||||||||
| Qwen3-max (w/ thinking) | 0.870.02 | 0.560.03 | 0.780.04 | 0.470.02 | 0.340.01 | 0.400.01 | 0.490.03 | 0.330.02 |
| Gemini-2.5-Flash (think on) | 0.580.01 | 0.310.04 | 0.500.02 | 0.290.03 | 0.140.01 | 0.200.04 | 0.270.04 | 0.220.02 |
| Doubao-Seed-1.6-Thinking | 0.630.03 | 0.400.01 | 0.610.03 | 0.310.02 | 0.280.04 | 0.310.01 | 0.270.03 | 0.210.02 |
| GLM-4.5 (w/ thinking) | 0.700.04 | 0.490.02 | 0.630.01 | 0.470.04 | 0.350.03 | 0.370.02 | 0.450.02 | 0.370.01 |
| GLM-4.6 (w/ thinking) | 0.750.01 | 0.560.04 | 0.680.02 | 0.490.01 | 0.400.04 | 0.400.03 | 0.470.04 | 0.360.02 |
| Claude-4-Sonnet (w/ thinking) | 0.690.03 | 0.460.01 | 0.580.04 | 0.430.02 | 0.360.01 | 0.360.04 | 0.390.03 | 0.310.02 |
| Claude-4.5-Sonnet (w/ thinking) | 0.730.02 | 0.570.00 | 0.650.02 | 0.510.03 | 0.410.02 | 0.410.01 | 0.350.03 | 0.290.02 |
| deepseek-R1-0528 (w/ thinking) | 0.530.01 | 0.370.01 | 0.460.02 | 0.420.04 | 0.390.03 | 0.360.02 | 0.330.03 | 0.260.02 |
| gpt-5.2 (w/ thinking) | 0.810.04 | 0.610.03 | 0.720.03 | 0.530.01 | 0.450.04 | 0.450.03 | 0.490.03 | 0.370.03 |
| o3 (w/ thinking) | 0.700.02 | 0.570.01 | 0.680.02 | 0.500.02 | 0.460.01 | 0.440.04 | 0.460.03 | 0.350.02 |
| Gemini-2.5-Pro | 0.600.02 | 0.370.03 | 0.550.01 | 0.490.04 | 0.340.02 | 0.380.03 | 0.370.03 | 0.300.03 |
3 Experiments
3.1 Experimental Setup
Model Zoo. We evaluate a diverse set of 24 representative large language models covering different architectures and parameter scales. Our evaluation benchmark encompasses a wide spectrum of proprietary models, including the OpenAI o-series (o3) (OpenAI, 2025e) and GPT series (GPT-4.1 (OpenAI, 2025a), GPT-5.2 (OpenAI, 2025d, b, c)); the DeepSeek family, covering the R1 series (Guo et al., 2025) and V3 variants (V3-0324 (Liu et al., 2024), V3.1 (DeepSeekAI, 2025), V3.2 (Liu et al., 2025)); Anthropic’s Claude series (Claude-4-Sonnet (Anthropic, 2025a), Claude-4.5-Sonnet (Anthropic, 2025b)); and Google’s Gemini-2.5 suite (Flash and Pro (Comanici et al., 2025; Google, 2025b, a)). Additionally, we include other leading models such as Qwen3-Max (Team, 2025), GLM variants (GLM-4.5 (Zeng et al., 2025b), GLM-4.6 (z.ai, 2025)), Seed-1.6 (ByteDance, 2025), and LongCat-Flash (Team et al., 2025d). To ensure a fair comparison, we distinguish between reasoning-enhanced and non reasoning models. Hybrid architectures capable of toggling between modes are assessed in both states independently.
Datasets. To systematically evaluate agent robustness in complex environments, we delineate core capabilities into two dimensions: (1) static knowledge retrieval and (2) dynamic multi-turn interaction. Specifically, for static retrieval tasks, we employ two representative benchmarks: 2WikiMultiHopQA (Ho et al., 2020) and HotpotQA (Yang et al., 2018). Collectively, the averaged performance across these two benchmarks is reported as our ’Search’ metric, representing static knowledge retrieval. Given that these tasks inherently lack user interaction and solely involve tool calls, we primarily introduce tool noise. Conversely, for dynamic interaction tasks, we utilize VitaBench and -Bench, two benchmarks that support the simultaneous injection of both user noise and tool noise. In each dimension, noisy evaluation instances are constructed by perturbing the original benchmark tasks as base datasets. Across all test domains, noise is introduced by perturbing key interaction interfaces, while strictly preserving the invariance of the original task objectives. Finally, to strike a balance between the experimental necessity of covering all predefined noise types and the practical constraints of evaluation scalability and cost, we adopt a stratified sampling strategy: 25% of samples are drawn from each scenario within VitaBench and -Bench, and 500 instances are sampled from the multi-hop QA datasets. This design ensures precise comparative analysis of model performance in both noisy and noise-free environments under a unified task distribution.
Metrics. We report mean performance over independent trials per task. Our evaluation captures both effectiveness and efficiency: Avg@4: The average task success rate across trials, serving as the primary measure of agent robustness (effectiveness). Avg_Tokens@4: The average token consumption across trials, quantifying computational cos Avg_Steps@4: The average number of interaction steps required to complete the task across trials, quantifying computational cost. Robustness: The relative rate of change of a metric above from noise-free to noisy environments, quantifying agent robustness under noisy conditions.
3.2 Noise Robustness Remains an Open Challenge for Current Agents
To assess whether noise robustness constitutes a critical bottleneck for agent deployment, we evaluated a series of state-of-the-art models following the experimental settings described in Section 3.1. Table 1 summarizes the performance of each model under both non-noise and noisy environments. Based on the experimental results, we make the following observation:
Observation 1: Agents universally exhibit significant vulnerability to noise, with varying degrees of sensitivity. Experimental result reveals that nearly all models suffer severe performance degradation upon the injection of noise, with an average accuracy drop of up to 20.8%. This performance loss is prevalent across all evaluation domains; for instance, the average declines in the Vitbench, -bench, and Search domains reach 24.1%, 21.2%, and 16.8%, respectively. Furthermore, we observe significant disparities in the magnitude of performance degradation across different models when subjected to identical noise interference. These results not only reveal the significant performance gap between an ideal noise-free environment and a real noisy scenario, but also highlight the intrinsic differences in models’ sensitivity to noise, posing a serious challenge to the widespread deployment of current agents.
3.3 Misalignment Exists between Reasoning Ability and Robustness
To further explore the relationship between general reasoning ability and noise robustness, we separately calculated the robustness scores of thinking models and non-thinking models. The comparative results are visualized in Figure 3, from which we summarize the following finding:
Observation 2: Reasoning ability and robustness are not closely related. Our experimental results indicate that reasoning ability and noise robustness are not closely related. Specifically, as shown in Figure 3, with the exception of specific scenarios (e.g., Delivery), the robustness scores of thinking models are relatively lower than those of non-thinking models across most other scenarios. This performance disadvantage prevalent across multiple domains clearly indicates that strong reasoning ability do not inherently confer robustness against noise. To deeply investigate the root causes of the low robustness observed in powerful reasoning models under noisy conditions, we analyzed the models’ reasoning trajectorys. Our study uncovered an interesting phenomenon: reasoning models often fail to effectively filter out noisy information. Instead, they tend to misinterpret this noise as meaningful, critical signals, and on this basis, attempt to construct complex logical chains to rationalize the noise. This mechanism leads to the generation of spurious reasoning chains—where the model produces solutions that appear logically rigorous but are in fact incorrect. The fundamental reason for this behavior is closely related to the training of reasoning models; their training data is typically strictly screened and pre-processed to eliminate noise. Consequently, when noise appears during the testing phase, the powerful reasoning capability of these models more readily fabricates information to make the noise seem rational.
3.4 How Noise Disrupts Reasoning via Entropy Injection
To investigate the microscopic mechanisms of how noise disrupts task execution, we conducted an analysis of step-wise entropy. We sampled 50 trajectories across 6 scenarios, comparing step-wise entropy evolution under three conditions: Non-Noise, User-Noise, and Tool-Noise. Figure 4 illustrates these dynamics. We make the following observations:
Observation 3: Noise alter entropy via distinct mechanisms. The results indicate that different noise sources alter entropy changes in the reasoning process through fundamentally different mechanisms. In noise-free baselines, we observe entropy decay, where the agent progressively resolves uncertainty, converging towards a deterministic solution. In contrast, user noise induces a high intercept entropy profile, characterized by extreme uncertainty at the initial step. This reflects the cognitive load required to disambiguate user intent before any planning can commence. Crucially, as shown in Figure 4, tool noise results in a relatively high level of entropy starting from the middle stage and persisting through the late stage of the agent trajectory. Here, the output of hallucination-inducing tools introduces uncertain information into the context, making it difficult for the agent to establish a reliable chain of reasoning.
3.5 Significant Performance Disparities Induced by Noise from Different Sources and Granularities
To comprehensively characterize the distinct impact of noise on agent performance, we conduct a two-level analysis from both the macroscopic scenario and the microscopic noise granularity, as shown in Figure 5(b) and Figure 5(a), respectively. Based on the experimental results, we summarize the following observations:
Observation 4: Fine-grained noise types exhibit markedly different destructive effects on agent performance. Our results demonstrate that the impact of noise is highly non-uniform and strongly depends on the specific noise type.
On the tool side, execution failures are the most destructive, reducing the average performance score to 0.31. This severe degradation arises because such noise induces structural execution breakdowns, which sever the agent’s access to environmental feedback. As a result, the reasoning process can no longer progress due to the absence of factual grounding. In contrast, noise types such as incomplete information or redundancy exhibit relatively milder effects. This suggests that agents retain a limited ability to filter and complete partial information, enabling them to extract useful signals from noisy tool outputs and sustain basic decision-making chains.
On the user side, instruction conflicts are the most harmful. When confronted with logically contradictory instructions, agents struggle to infer a consistent and correct intent, leading to pronounced performance degradation. By contrast, instruction ambiguity causes only minor performance drops. This indicates that although ambiguous intent increases interpretation difficulty, agents can often rely on the strong prior knowledge embedded in large language models to perform probabilistic semantic repair, thereby partially mitigating the adverse impact of noise.
Observation 5: Agents exhibit substantially higher sensitivity to tool-side noise than to user-side noise. Empirical results reveal a pronounced robustness gap between the two noise sources: across all agent families, vulnerability to tool-side noise is significantly higher than to user-side noise. Quantitatively, tool noise typically leads to severe performance drops of 20%–30%, whereas user noise generally results in reductions of around 10
To explain this asymmetric sensitivity, we further analyze the agent–environment interaction process and identify a fundamental distinction between user noise and environmental (tool) noise during reasoning.
User noise is semantically repairable. User noise primarily manifests as perturbations at the linguistic level. During pretraining, large language models acquire extensive linguistic and world knowledge, endowing them with strong error-correction and associative capabilities. Consequently, even when user instructions are noisy, agents can often internally repair these perturbations at the semantic level, allowing task execution to proceed.
Tool noise causes a breakdown of factual grounding. In contrast, tool-side noise directly corrupts the agent’s only channel for perceiving the external world. During task execution, tool outputs serve as the sole objective evidence for state estimation and action selection. Once this factual grounding is disrupted, the loss is structural rather than semantic: agents cannot infer the true environment state through internal reasoning alone. As a result, the reasoning chain loses its support and collapses, ultimately leading to task failure.
3.6 When Does Noise Injection Do the Most Impact
To investigate how the timing of noise injection affects agent performance, we consider three injection stages: early, middle, and late. We conduct experiments on four representative model families, including Claude, Gemini, DeepSeek, and LongCat. The results are shown in Figure 6, with additional results for more models reported in Appendix A.2. Based on these results, we derive the following observation:
Observation 6: Agents are most sensitive to noise injected at the middle stage. As illustrated in Figure 6, across all four models, injecting noise during the middle stage leads to a significantly larger performance degradation than injecting noise at either the early or late stage. This indicates that agents exhibit relatively higher robustness to noise introduced at the early or the late of the trajectory, while being more sensitive to noise during the middle stage. Importantly, the consistency of this trend across models with different architectures and reasoning strategies suggests that the heightened sensitivity during the intermediate stage of the process may be an inherent property, rather than a model-specific phenomenon.
4 Related Work
4.1 Benchmarks for Tool-Using and Interactive Agents
Recent advances have significantly expanded the scope of agent evaluation from single-turn tasks (Huang et al., 2023; Qin et al., 2023; Patil et al., ) to complex reasoning scenarios involving tool use and multi-step decision making (Farn and Shin, 2023; Wang et al., 2023; Qian et al., 2024). A growing body of work focuses on agents that iteratively interact with external tools to solve problems (Lu et al., 2025; Wang et al., 2025c; Qian et al., 2025), establishing protocols to assess planning capabilities (Yao et al., 2022; Schick et al., 2023; Zhou et al., 2023). Furthermore, benchmarks like -Bench (Barres et al., 2025) and VitaBench (He et al., 2025) have been developed to evaluate agents in multi-turn interactions where maintaining state is crucial.
However, a significant gap remains between benchmark environments and real-world applications. Most existing frameworks rely on idealized assumptions (Yang et al., 2024; Wang et al., 2025d), providing agents with carefully curated instructions and well-controlled environments. Unlike prior work, AgentNoiseBench introduces an automated pipeline capable of injecting noise into any existing agent-centric benchmarks while maintaining task solvability.
4.2 Robustness and Degradation of Agents under Noise
The deployment of agents in complex real-world environments has highlighted the critical importance of robustness (Li et al., 2025b; Levy et al., 2025; Agrawal et al., 2025). Recent studies indicate that agents suffer substantial performance degradation when faced with out-of-distribution scenarios or ambiguous user behaviors (Anghel et al., 2025; Wan et al., 2025; Herrera-Poyatos et al., 2025; Wang et al., 2025b). Various efforts have explored agent behavior under prompt perturbations or dialog variations (Li et al., 2025a; Deshpande et al., 2025; Qi et al., 2025). On the execution side, reliance on external tools introduces further instability, where agents usually encounter noisy, incomplete, or failed outputs (Xu et al., 2024b; Zhang et al., 2024b; Vuddanti et al., 2025), often leading to compounding errors (Yang et al., 2025; Zhu et al., 2025).
Nevertheless, existing evaluations (Nalbandyan et al., 2025; Wen et al., 2025; Yu et al., 2025) generally lack a systematic noise taxonomy and often focus on single noise sources. In addition, they typically only use the success rate as an evaluation metric. AgentNoiseBench introduces a unified framework covering both instruction and tool-execution noise and employs a trajectory-aware evaluation protocol to provide a multi-dimensional analysis to comprehensively characterize agent performance under noise.
5 Limitations
Despite the insights of our benchmark, several limitations remain. First, the evaluation targets language-centric, tool-augmented agents and may not generalize to agents with explicit planning, learned world models, or asynchronous tool environments. Second, while trajectory-level entropy is a useful diagnostic signal, it provides limited causal insight, as the analysis captures correlations between noise and instability. Third, trajectory-aware evaluation incurs computational overhead, and the benchmark lacks a scalable trade-off between diagnostic fidelity and cost, limiting its applicability in large-scale or real-world settings.
6 Future Work
These limitations motivate future research. First, we will study robustness-oriented training, such as reinforcement learning with noise-aware objectives, to improve agent stability under noisy interactions. Second, we plan to extend the benchmark to a wider range of agent architectures, task domains, and tool ecosystems. Finally, future work will develop analytical frameworks integrating trajectory-level diagnostics with causal and counterfactual analysis to better understand agent robustness.
7 Conclusion
In this work, we rethink the evaluation of LLM-based agents through the lens of real-world imperfections, introducing AgentNoiseBench to bridge the gap between controlled benchmarks and practical deployments. With our automated pipeline, we can explicitly incorporate real-world noise into any agent-centric benchmarks while ensuring the tasks are still solvable. Our evaluation reveals that nearly all models suffer substantial performance degradation, with an average accuracy drop of 20.8%. We urge greater attention to real-world imperfections when training and evaluating agents.
Appendix A More Experimental Results
A.1 Main Results
In this section, we conduct a more fine-grained analysis of the Agent Noise Benchmark results reported in Table 1. Specifically, we investigate how different types of user-side noise and tool-side noise impact mainstream agentic models across varied task scenarios.
Our analysis focuses on examining the granular effects of individual noise types on model performance. We systematically evaluate how each specific noise category—ranging from user ambiguity and topic shifts to tool errors and incomplete outputs—affects agent behavior across different application domains (airline, delivery, in-store, and retail). This fine-grained perspective allows us to identify which noise types pose the greatest challenges to current agentic systems and how these vulnerabilities vary across different task contexts and model architectures.
Detailed results are presented in Table 2, Table 3, Table 4, Table 5, Table 6, Table 7, Table 8, and Table 9. All experimental settings follow those described in Section 3.
Results show that all evaluated agents experience performance degradation under at least one noise type, with sensitivity patterns depending on both noise origin and model architecture. User-side noise primarily affects intent understanding and planning stages, leading to instruction misinterpretation or suboptimal plans. Tool-side noise more directly impairs execution reliability by introducing uncertainty in tool outputs. Notably, certain noise types such as conflict and incomplete information consistently cause more severe performance drops across models, while others like redundancy show more varied impacts depending on the specific agent architecture.
These findings emphasize the importance of systematically assessing and enhancing agent robustness in noisy environments at a granular level, and highlight the need for noise-type-specific mitigation strategies tailored to address the unique challenges posed by different categories of user-side and tool-side disturbances.
| Model | Airline | |||||
| clean | ambiguity | topic_shift | conflict | redundancy | boundary | |
| Non-think Models | ||||||
| DeepSeek-V3-0324 | 50.02.3 | 45.81.9 | 50.02.1 | 50.01.7 | 45.82.4 | 50.01.8 |
| Doubao-Seed-1.6 | 62.52.2 | 54.21.8 | 54.22.0 | 54.22.5 | 54.22.3 | 62.51.9 |
| GPT-4.1 | 60.42.1 | 41.72.6 | 47.91.8 | 47.92.2 | 41.72.7 | 60.42.0 |
| DeepSeek-V3.1 | 68.81.9 | 54.22.4 | 56.32.2 | 52.12.6 | 54.21.8 | 68.82.5 |
| gemini-2.5-pro | 77.12.2 | 54.22.1 | 60.42.3 | 62.51.9 | 62.52.0 | 77.12.4 |
| Qwen3-Max | 66.72.4 | 41.72.2 | 54.22.5 | 54.21.8 | 54.22.1 | 66.72.3 |
| LongCat-Flash-Chat | 66.72.0 | 50.02.3 | 58.32.2 | 56.31.9 | 56.32.4 | 66.72.1 |
| Claude-4-Sonnet | 77.12.5 | 66.72.0 | 60.42.4 | 58.32.1 | 66.72.3 | 77.11.9 |
| Claude-4.5-Sonnet | 75.02.1 | 68.82.5 | 60.41.9 | 58.32.2 | 60.42.6 | 75.02.3 |
| Thinking Models | ||||||
| Qwen3-max (w/ thinking) | 89.62.4 | 70.82.2 | 58.32.6 | 58.32.0 | 52.12.3 | 89.62.5 |
| Gemini-2.5-Flash (think on) | 77.12.3 | 56.31.9 | 54.22.2 | 54.22.4 | 56.32.1 | 77.12.6 |
| Doubao-Seed-1.6-Thinking | 77.12.5 | 54.22.1 | 60.42.3 | 60.41.9 | 56.32.4 | 77.12.0 |
| GLM-4.6 (w/ thinking) | 75.02.3 | 77.12.4 | 66.72.0 | 66.72.2 | 68.82.5 | 75.02.1 |
| deepseek-R1-0528 (w/ thinking) | 70.82.4 | 52.12.6 | 52.12.2 | 52.12.5 | 41.72.0 | 70.82.3 |
| gpt-5.2 (w/ thinking) | 79.22.5 | 60.42.3 | 68.82.1 | 66.72.4 | 66.72.6 | 79.22.2 |
| o3 (w/ thinking) | 77.12.2 | 72.92.4 | 66.72.3 | 66.72.1 | 64.62.5 | 77.12.6 |
| Model | Delivery | |||||
| clean | ambiguity | topic_shift | conflict | redundancy | boundary | |
| Non-think Models | ||||||
| DeepSeek-V3-0324 | 42.32.6 | 33.01.9 | 38.02.4 | 34.01.7 | 38.02.5 | 49.02.1 |
| Doubao-Seed-1.6 | 35.02.4 | 27.01.8 | 33.02.2 | 32.02.5 | 37.02.3 | 38.01.9 |
| GPT-4.1 | 38.02.1 | 33.02.6 | 40.01.8 | 35.02.4 | 38.01.9 | 38.02.5 |
| DeepSeek-V3.1 | 33.02.5 | 22.02.0 | 32.02.4 | 38.01.8 | 39.02.6 | 42.62.3 |
| gemini-2.5-pro | 43.02.3 | 44.02.1 | 46.02.5 | 48.02.2 | 48.01.9 | 50.02.4 |
| Qwen3-Max | 27.02.6 | 27.02.1 | 25.02.4 | 24.02.3 | 31.02.0 | 31.01.8 |
| LongCat-Flash-Chat | 35.02.3 | 42.01.9 | 44.02.1 | 41.02.5 | 42.02.4 | 44.02.0 |
| Claude-4-Sonnet | 35.02.4 | 29.02.2 | 42.02.6 | 39.01.8 | 40.02.5 | 45.02.1 |
| Claude-4.5-Sonnet | 47.02.0 | 48.02.5 | 53.02.3 | 47.02.6 | 54.02.1 | 53.61.9 |
| Thinking Models | ||||||
| DeepSeek-R1-0528 (w/ thinking) | 41.02.6 | 35.02.4 | 48.02.5 | 41.02.0 | 50.02.3 | 44.02.2 |
| Doubao-Seed-1.6-Thinking | 35.02.5 | 35.02.0 | 41.02.4 | 32.02.6 | 37.02.3 | 37.02.1 |
| Gemini-2.5-Flash (think on) | 26.02.4 | 18.02.1 | 35.02.3 | 29.02.5 | 21.02.6 | 34.01.9 |
| gemini-2.5-pro (w/ thinking) | 43.02.3 | 44.02.7 | 46.02.4 | 48.02.1 | 48.02.5 | 50.02.2 |
| GLM-4.5 (w/ thinking) | 37.02.4 | 38.02.6 | 45.02.2 | 43.02.3 | 39.62.1 | 46.02.5 |
| GLM-4.6 (w/ thinking) | 38.62.3 | 29.02.6 | 47.52.4 | 45.22.2 | 41.42.5 | 47.02.1 |
| GPT-5.2 (w/ thinking) | 51.02.5 | 53.02.3 | 53.02.6 | 48.02.4 | 51.02.2 | 54.12.7 |
| o3 (w/ thinking) | 39.02.3 | 39.42.6 | 55.02.4 | 47.02.2 | 56.02.5 | 46.02.1 |
| Model | Instore | |||||
| clean | ambiguity | topic_shift | conflict | redundancy | boundary | |
| Non-think Models | ||||||
| DeepSeek-V3-0324 | 30.02.5 | 18.61.8 | 25.02.3 | 16.01.9 | 22.02.6 | 24.02.1 |
| Doubao-Seed-1.6 | 53.92.4 | 38.21.7 | 48.22.2 | 35.62.5 | 42.32.3 | 50.21.8 |
| GPT-4.1 | 59.02.1 | 42.02.6 | 52.01.8 | 40.02.4 | 48.01.9 | 54.02.5 |
| DeepSeek-V3.1 | 55.02.6 | 39.62.0 | 49.42.4 | 37.61.7 | 45.52.5 | 51.62.3 |
| gemini-2.5-pro | 53.02.2 | 37.62.3 | 47.12.5 | 38.02.1 | 44.01.9 | 49.02.4 |
| Qwen3-Max | 57.82.6 | 40.02.1 | 50.82.4 | 38.02.3 | 46.02.0 | 52.61.8 |
| LongCat-Flash-Chat | 62.02.3 | 43.21.9 | 54.02.1 | 42.02.5 | 50.02.4 | 57.02.0 |
| Claude-4-Sonnet | 56.02.4 | 39.02.2 | 50.02.6 | 38.01.8 | 46.02.5 | 52.02.1 |
| Claude-4.5-Sonnet | 58.02.0 | 41.62.5 | 52.02.3 | 40.02.6 | 48.02.1 | 54.01.9 |
| Thinking Models | ||||||
| DeepSeek-R1-0528 (w/ thinking) | 59.02.6 | 42.72.4 | 53.02.5 | 41.02.0 | 49.02.3 | 55.02.2 |
| Doubao-Seed-1.6-Thinking | 45.02.5 | 32.52.0 | 40.02.4 | 30.02.6 | 37.02.3 | 42.02.1 |
| Gemini-2.5-Flash (think on) | 26.02.4 | 16.02.1 | 20.02.3 | 14.02.5 | 18.02.6 | 22.01.9 |
| gemini-2.5-pro (w/ thinking) | 57.02.3 | 40.62.7 | 51.02.4 | 39.02.1 | 47.02.5 | 53.02.2 |
| GLM-4.5 (w/ thinking) | 64.02.4 | 46.02.6 | 57.02.2 | 44.02.3 | 52.02.1 | 59.02.5 |
| GLM-4.6 (w/ thinking) | 65.82.3 | 48.52.6 | 59.42.4 | 46.02.2 | 54.02.5 | 61.02.1 |
| GPT-5.2 (w/ thinking) | 65.02.5 | 47.92.3 | 58.02.6 | 45.02.4 | 53.02.2 | 60.02.7 |
| o3 (w/ thinking) | 64.02.3 | 47.52.6 | 58.02.4 | 46.02.2 | 54.02.5 | 60.02.1 |
| Model | Retail | |||||
| clean | ambiguity | topic_shift | conflict | redundancy | boundary | |
| Non-think Models | ||||||
| DeepSeek-V3-0324 | 66.12.4 | 47.81.8 | 55.32.3 | 51.71.9 | 57.62.5 | 59.42.1 |
| Doubao-Seed-1.6 | 70.52.3 | 51.81.7 | 59.92.2 | 56.22.4 | 62.12.1 | 63.81.8 |
| GPT-4.1 | 74.12.1 | 56.22.6 | 64.11.8 | 60.02.5 | 65.61.9 | 68.22.3 |
| DeepSeek-V3.1 | 72.32.5 | 53.72.0 | 62.22.4 | 58.11.7 | 64.12.6 | 66.12.3 |
| gemini-2.5-pro | 66.12.2 | 48.52.3 | 56.92.5 | 52.32.1 | 58.41.9 | 60.42.4 |
| Qwen3-Max | 71.42.6 | 52.72.1 | 61.02.4 | 56.42.3 | 62.62.0 | 65.01.8 |
| LongCat-Flash-Chat | 79.52.3 | 59.61.9 | 68.22.1 | 63.42.5 | 69.52.4 | 71.82.0 |
| Claude-4-Sonnet | 73.22.4 | 54.62.2 | 62.82.6 | 59.21.8 | 64.32.5 | 66.92.1 |
| Claude-4.5-Sonnet | 79.52.0 | 59.82.5 | 68.62.3 | 63.62.6 | 69.42.1 | 72.21.9 |
| Thinking Models | ||||||
| DeepSeek-R1-0528 (w/ thinking) | 57.12.6 | 41.72.4 | 49.32.5 | 45.92.0 | 51.92.3 | 53.42.2 |
| Doubao-Seed-1.6-Thinking | 58.02.5 | 42.42.0 | 50.12.4 | 46.82.6 | 53.12.3 | 55.02.1 |
| Gemini-2.5-Flash (think on) | 69.62.4 | 51.12.1 | 59.32.3 | 55.32.5 | 61.62.6 | 63.61.9 |
| gemini-2.5-pro (w/ thinking) | 73.22.3 | 55.22.7 | 63.12.4 | 59.12.1 | 64.82.5 | 66.92.2 |
| GLM-4.5 (w/ thinking) | 80.42.4 | 62.12.6 | 70.32.2 | 65.82.3 | 71.52.1 | 74.02.5 |
| GLM-4.6 (w/ thinking) | 84.82.3 | 65.32.6 | 73.82.4 | 68.32.2 | 74.32.5 | 76.52.1 |
| GPT-5.2 (w/ thinking) | 73.22.5 | 55.22.3 | 63.12.6 | 59.12.4 | 64.82.2 | 66.82.7 |
| o3 (w/ thinking) | 80.02.3 | 62.22.6 | 70.92.4 | 67.02.2 | 72.92.5 | 75.12.1 |
| Model | Airline | |||||
| clean | error | failure | incomplete | induce | redundancy | |
| Non-think Models | ||||||
| DeepSeek-V3-0324 | 50.02.0 | 48.52.2 | 50.01.6 | 50.02.1 | 45.82.0 | 45.82.3 |
| Doubao-Seed-1.6 | 62.51.9 | 60.92.1 | 54.21.5 | 47.92.0 | 54.21.8 | 54.22.1 |
| GPT-4.1 | 60.42.4 | 61.22.5 | 47.92.0 | 43.82.3 | 41.82.2 | 47.92.4 |
| DeepSeek-V3.1 | 68.82.1 | 68.32.3 | 52.11.8 | 56.32.2 | 54.22.1 | 54.22.2 |
| gemini-2.5-pro | 77.12.5 | 75.92.4 | 60.41.9 | 52.12.3 | 54.22.4 | 62.52.5 |
| Qwen3-Max | 66.72.3 | 65.12.4 | 54.21.9 | 45.82.2 | 41.72.3 | 54.22.4 |
| LongCat-Flash-Chat | 66.72.1 | 67.62.2 | 58.31.7 | 54.22.1 | 50.02.0 | 56.32.2 |
| Claude-4-Sonnet | 77.12.2 | 74.72.3 | 60.41.9 | 58.32.2 | 66.72.1 | 66.72.3 |
| Claude-4.5-Sonnet | 75.02.4 | 78.12.4 | 58.32.0 | 60.42.3 | 68.82.3 | 60.42.5 |
| Thinking Models | ||||||
| Qwen3-max (w/ thinking) | 89.62.6 | 89.12.5 | 58.32.1 | 60.42.5 | 70.82.6 | 52.12.7 |
| Gemini-2.5-Flash (think on) | 77.12.0 | 76.62.2 | 54.21.7 | 47.92.1 | 56.32.0 | 56.32.2 |
| Doubao-Seed-1.6-Thinking | 77.12.1 | 76.62.3 | 60.41.8 | 60.42.2 | 54.22.1 | 56.32.3 |
| GLM-4.6 (w/ thinking) | 77.12.6 | 74.52.5 | 66.72.1 | 62.52.5 | 77.12.6 | 68.82.7 |
| deepseek-R1-0528 (w/ thinking) | 70.82.2 | 70.32.3 | 52.11.9 | 50.02.3 | 52.12.2 | 41.72.4 |
| gpt-5.2 (w/ thinking) | 79.22.8 | 78.72.7 | 66.72.3 | 68.82.7 | 60.42.8 | 66.72.9 |
| o3 (w/ thinking) | 77.12.5 | 76.62.5 | 66.72.1 | 54.22.5 | 72.92.6 | 64.62.7 |
| Model | Delivery | |||||
| clean | error | failure | incomplete | induce | redundancy | |
| Non-think Models | ||||||
| DeepSeek-V3-0324 | 42.32.4 | 49.01.8 | 16.82.2 | 37.21.9 | 39.82.6 | 40.52.1 |
| Doubao-Seed-1.6 | 39.82.3 | 38.31.7 | 14.22.0 | 34.62.4 | 36.32.2 | 37.91.8 |
| GPT-4.1 | 48.62.1 | 38.02.6 | 22.11.8 | 42.92.5 | 45.71.9 | 46.32.3 |
| DeepSeek-V3.1 | 44.22.5 | 42.62.0 | 18.52.3 | 39.41.7 | 41.62.4 | 42.82.6 |
| gemini-2.5-pro | 51.32.2 | 50.02.3 | 21.32.5 | 46.72.1 | 49.21.9 | 50.12.4 |
| Qwen3-Max | 47.52.6 | 31.02.1 | 19.82.4 | 41.32.3 | 44.62.0 | 45.21.8 |
| LongCat-Flash-Chat | 40.72.3 | 44.01.9 | 15.92.1 | 35.82.5 | 38.22.4 | 39.42.0 |
| Claude-4-Sonnet | 45.92.4 | 45.02.2 | 20.32.6 | 40.81.8 | 43.12.5 | 44.62.1 |
| Claude-4.5-Sonnet | 49.32.0 | 53.62.5 | 22.52.3 | 44.22.6 | 46.82.1 | 47.91.9 |
| Thinking Models | ||||||
| Qwen3-max (w/ thinking) | 58.72.3 | 44.32.7 | 25.62.4 | 52.82.1 | 55.92.5 | 56.32.2 |
| Gemini-2.5-Flash (think on) | 41.22.4 | 29.82.1 | 16.42.3 | 36.72.5 | 38.92.6 | 39.71.9 |
| Doubao-Seed-1.6-Thinking | 43.92.5 | 32.72.0 | 18.92.4 | 39.42.6 | 41.32.3 | 42.12.1 |
| GLM-4.6 (w/ thinking) | 54.82.4 | 42.32.6 | 24.92.2 | 50.12.3 | 52.72.1 | 53.22.5 |
| deepseek-R1-0528 (w/ thinking) | 46.32.6 | 35.92.4 | 21.42.5 | 42.12.0 | 44.72.3 | 45.32.2 |
| gpt-5.2 (w/ thinking) | 62.12.5 | 48.62.3 | 28.72.6 | 56.32.4 | 59.42.2 | 60.22.7 |
| o3 (w/ thinking) | 53.72.3 | 41.52.6 | 24.22.4 | 49.32.2 | 51.82.5 | 52.52.1 |
| Model | Instore | |||||
| clean | error | failure | incomplete | induce | redundancy | |
| Non-think Models | ||||||
| DeepSeek-V3-0324 | 30.02.4 | 29.01.8 | 24.02.3 | 8.01.9 | 28.02.5 | 30.02.1 |
| Doubao-Seed-1.6 | 53.92.3 | 46.51.7 | 33.02.2 | 35.72.4 | 45.02.1 | 53.91.8 |
| GPT-4.1 | 59.02.1 | 56.02.6 | 41.01.8 | 13.02.5 | 48.01.9 | 59.02.3 |
| DeepSeek-V3.1 | 55.02.5 | 59.02.0 | 36.02.4 | 30.01.7 | 49.02.6 | 55.02.3 |
| gemini-2.5-pro | 57.02.2 | 55.02.3 | 34.02.5 | 24.02.1 | 46.01.9 | 57.02.4 |
| Qwen3-Max | 57.82.6 | 50.02.1 | 27.02.4 | 11.02.3 | 44.02.0 | 57.81.8 |
| LongCat-Flash-Chat | 62.02.3 | 63.01.9 | 38.02.1 | 29.02.5 | 49.02.4 | 62.02.0 |
| Claude-4-Sonnet | 58.32.4 | 53.92.2 | 45.52.6 | 25.21.8 | 46.42.5 | 58.32.1 |
| Claude-4.5-Sonnet | 61.32.0 | 55.82.5 | 44.02.3 | 41.62.6 | 55.02.1 | 61.31.9 |
| Thinking Models | ||||||
| Qwen3-max (w/ thinking) | 57.02.3 | 55.02.7 | 34.02.4 | 24.02.1 | 46.02.5 | 57.02.2 |
| DeepSeek-R1-0528 (w/ thinking) | 59.02.6 | 56.02.4 | 46.02.5 | 25.02.0 | 46.02.3 | 59.02.2 |
| Doubao-Seed-1.6-Thinking | 45.02.5 | 45.02.0 | 36.02.4 | 36.02.6 | 36.02.3 | 45.02.1 |
| Gemini-2.5-Flash (think on) | 26.02.4 | 24.02.1 | 10.02.3 | 2.02.5 | 18.02.6 | 26.01.9 |
| GLM-4.6 (w/ thinking) | 65.82.4 | 59.02.6 | 57.02.2 | 53.52.3 | 45.02.1 | 65.82.5 |
| GPT-5.2 (w/ thinking) | 65.02.5 | 57.02.3 | 54.02.6 | 42.02.4 | 49.02.2 | 65.02.7 |
| o3 (w/ thinking) | 64.02.3 | 59.02.6 | 53.02.4 | 62.02.2 | 58.02.5 | 64.02.1 |
| Model | Retail | |||||
| clean | error | failure | incomplete | induce | redundancy | |
| Non-think Models | ||||||
| DeepSeek-V3-0324 | 66.12.4 | 38.51.8 | 24.12.3 | 29.51.9 | 42.32.5 | 58.22.1 |
| Doubao-Seed-1.6 | 70.52.3 | 45.21.7 | 28.62.2 | 35.62.4 | 48.92.1 | 62.41.8 |
| GPT-4.1 | 74.12.1 | 52.92.6 | 26.81.8 | 42.72.5 | 55.31.9 | 66.82.3 |
| DeepSeek-V3.1 | 72.32.5 | 48.42.0 | 33.92.4 | 38.81.7 | 51.82.6 | 64.52.3 |
| gemini-2.5-pro | 66.12.2 | 41.22.3 | 28.62.5 | 34.02.1 | 47.11.9 | 59.32.4 |
| Qwen3-Max | 71.42.6 | 46.42.1 | 27.72.4 | 37.12.3 | 49.32.0 | 63.71.8 |
| LongCat-Flash-Chat | 79.52.3 | 54.31.9 | 33.92.1 | 43.12.5 | 57.22.4 | 71.82.0 |
| Claude-4-Sonnet | 73.22.4 | 49.32.2 | 29.52.6 | 39.91.8 | 52.02.5 | 65.62.1 |
| Claude-4.5-Sonnet | 79.52.0 | 55.52.5 | 39.32.3 | 45.32.6 | 58.12.1 | 72.41.9 |
| Thinking Models | ||||||
| Qwen3-Max (w/ thinking) | 75.92.3 | 51.12.7 | 40.22.4 | 42.52.1 | 56.52.5 | 68.92.2 |
| Gemini-2.5-Flash (think on) | 69.62.4 | 44.82.1 | 25.02.3 | 36.02.5 | 50.32.6 | 62.11.9 |
| Doubao-Seed-1.6-Thinking | 58.02.5 | 35.12.0 | 20.52.4 | 28.82.6 | 41.72.3 | 52.42.1 |
| GLM-4.6 (w/ thinking) | 84.82.4 | 60.02.6 | 42.02.2 | 48.52.3 | 63.22.1 | 77.32.5 |
| DeepSeek-R1-0528 (w/ thinking) | 57.12.6 | 34.42.4 | 25.02.5 | 31.62.0 | 42.62.3 | 51.82.2 |
| GPT-5.2 (w/ thinking) | 73.22.5 | 49.92.3 | 34.82.6 | 40.82.4 | 54.52.2 | 66.72.7 |
| o3 (w/ thinking) | 80.02.3 | 56.02.6 | 48.02.4 | 47.02.2 | 60.02.5 | 73.52.1 |
| Model | HotpotQA | |||||
| clean | error | failure | incomplete | induce | redundancy | |
| Non-think models | ||||||
| DeepSeek-V3-0324 | 0.210.08 | 0.190.02 | 0.160.06 | 0.230.04 | 0.200.06 | 0.210.02 |
| Gemini-2.5-Flash (think off) | 0.290.04 | 0.280.06 | 0.120.02 | 0.250.02 | 0.230.02 | 0.240.06 |
| GPT-4.1 | 0.330.04 | 0.290.01 | 0.300.06 | 0.290.04 | 0.290.04 | 0.290.02 |
| DeepSeek-V3.1 | 0.400.06 | 0.310.06 | 0.290.02 | 0.380.04 | 0.370.04 | 0.380.04 |
| DeepSeek-V3.2-Exp (w/o thinking) | 0.430.02 | 0.370.02 | 0.320.06 | 0.420.04 | 0.410.02 | 0.430.06 |
| gpt-3.5-turbo (w/o thinking) | 0.180.04 | 0.140.03 | 0.130.01 | 0.160.02 | 0.170.06 | 0.130.04 |
| GLM-4.6 | 0.410.06 | 0.370.02 | 0.050.02 | 0.400.04 | 0.430.01 | 0.410.08 |
| LongCat-Flash-Chat | 0.360.04 | 0.260.02 | 0.220.01 | 0.340.02 | 0.320.01 | 0.320.06 |
| Claude-4-Sonnet | 0.340.04 | 0.280.08 | 0.240.04 | 0.340.04 | 0.350.02 | 0.330.01 |
| Claude-4.5-Sonnet | 0.350.03 | 0.320.02 | 0.230.06 | 0.330.04 | 0.360.06 | 0.350.06 |
| Thinking Models | ||||||
| Qwen3-max (w/ thinking) | 0.430.01 | 0.330.02 | 0.140.06 | 0.410.01 | 0.360.04 | 0.430.06 |
| Gemini-2.5-Flash (think on) | 0.270.04 | 0.210.04 | 0.130.02 | 0.250.06 | 0.230.08 | 0.260.08 |
| GLM-4.5 (w/ thinking) | 0.440.02 | 0.350.04 | 0.270.04 | 0.410.01 | 0.440.02 | 0.370.05 |
| GLM-4.6 (w/ thinking) | 0.450.04 | 0.340.02 | 0.260.03 | 0.390.01 | 0.390.08 | 0.400.06 |
| Model | Wiki | |||||
| clean | error | failure | incomplete | induce | redundancy | |
| Non-think models | ||||||
| DeepSeek-V3-0324 | 0.220.02 | 0.180.04 | 0.060.04 | 0.200.02 | 0.200.06 | 0.190.02 |
| Gemini-2.5-Flash (think off) | 0.270.04 | 0.200.06 | 0.030.08 | 0.240.04 | 0.230.06 | 0.250.06 |
| GPT-4.1 | 0.250.02 | 0.220.01 | 0.290.08 | 0.220.02 | 0.160.02 | 0.240.04 |
| DeepSeek-V3.1 | 0.440.02 | 0.420.03 | 0.300.02 | 0.440.04 | 0.470.06 | 0.470.04 |
| DeepSeek-V3.2-Exp (w/o thinking) | 0.440.02 | 0.420.03 | 0.300.02 | 0.440.04 | 0.470.06 | 0.470.04 |
| gpt-3.5-turbo (w/o thinking) | 0.440.02 | 0.420.03 | 0.300.02 | 0.440.04 | 0.470.06 | 0.470.04 |
| GLM-4.6 | 0.510.06 | 0.440.04 | 0.280.06 | 0.490.02 | 0.460.06 | 0.500.04 |
| LongCat-Flash-Chat | 0.400.01 | 0.320.02 | 0.160.07 | 0.370.02 | 0.390.04 | 0.390.02 |
| Claude-4-Sonnet | 0.400.02 | 0.320.03 | 0.100.04 | 0.380.03 | 0.380.02 | 0.310.03 |
| Claude-4.5-Sonnet | 0.320.04 | 0.290.02 | 0.090.06 | 0.310.02 | 0.340.02 | 0.310.04 |
| Thinking Models | ||||||
| Qwen3-max (w/ thinking) | 0.560.02 | 0.360.02 | 0.180.07 | 0.440.04 | 0.510.04 | 0.400.09 |
| Gemini-2.5-Flash (think on) | 0.270.06 | 0.180.04 | 0.080.04 | 0.210.06 | 0.190.06 | 0.280.04 |
| GLM-4.5 (w/ thinking) | 0.460.02 | 0.350.08 | 0.190.07 | 0.430.03 | 0.440.04 | 0.450.08 |
| GLM-4.6 (w/ thinking) | 0.470.08 | 0.330.02 | 0.200.02 | 0.390.01 | 0.430.02 | 0.410.02 |
A.2 Impact of Noise Injection Stage
We also evaluate the performance degradation of the remaining four model families under noise injection at different stages:
Appendix B Noise Case Study
To facilitate a clearer understanding of our experimental pipeline, we randomly sample evaluation instances from different agent tasks and environments, and construct noisy variants using the noise injection mechanisms defined in the Agent Noise Benchmark. These noisy examples are generated by perturbing either intermediate tool outputs, contextual information, or decision-relevant signals, depending on the specific noise configuration under study. For illustration purposes, we present representative examples in Example B (clean trajectory), Example B (noisy trajectory with mid-stage noise injection), and additional cases in subsequent examples. The following are examples of user-side noise:
The following are examples of tool-side noise:
Appendix C Agent Noise Robustness: Future Discussion
Although this study has preliminarily constructed an evaluation framework for agent noise robustness and revealed key phenomena, several critical paths remain to be explored to advance the field towards greater practicality and theoretical depth. The following outlines future research directions from four perspectives: refinement of noise modeling, innovation in training paradigms, expansion of the evaluation system, and theoretical understanding of robustness.
(1) From Categorical Noise to Continuous Spectrum and Composite Noise Modeling The current study categorizes noise into “user-side” and “tool-side” types, providing a clear initial analytical framework. However, real-world noise is far from binary. Future work should strive to construct a noise spectrum model with greater continuity and composability. For instance, user-side noise could extend from simple lexical perturbations (e.g., typos) to more complex phenomena like “intention drift” or “contextual contradictions in multi-turn dialogues.” Tool-side noise should consider “temporal correlation” (e.g., intermittent failures) and “state dependency” (e.g., errors triggered only under specific input conditions). A more significant challenge lies in simulating the combination and synergistic effects of multiple noise sources—scenarios where ambiguous user instructions encounter delayed and incomplete tool responses. Building and evaluating such composite scenarios will better approximate the complexity of real-world deployments.
(2) Paradigm Shift: From “Cleaned Training” to “Active Noise-Injected Training” Most existing agent training relies on meticulously curated instructions and stable environments, creating a “fundamental mismatch” between training and deployment contexts. Drawing inspiration from data augmentation techniques in computer vision and reinforcement learning that enhance model robustness, future agent training must systematically incorporate noise exposure. This goes beyond merely adding random perturbations to data; it necessitates developing curriculum learning strategies. Initially, controlled, manageable noise is injected during training. As the agent’s capabilities improve, the intensity, diversity, and unpredictability of the noise are gradually increased. Concurrently, adversarial training methods can be explored. By establishing a dynamic game between a “noise generator” and the agent, noise samples that effectively challenge the agent’s current weaknesses can be generated, thereby targetedly enhancing the robustness of its vulnerable points.
(3) From Static Benchmark Evaluation to Dynamic Simulation and Real User Interaction Evaluation AgentNoiseBench represents a significant step forward with static benchmarks, but the ultimate testing ground should be the dynamic, open real world. Future work must expand the evaluation system in two directions. First, developing high-fidelity, multimodal environment simulators, such as GUI operation simulators incorporating random network latency, rendering errors of interface elements, or unstructured document outputs, would make the injection of tool-execution noise more ecologically valid. Second, exploring online evaluation based on real user interactions is crucial. Through crowdsourcing platforms or pilot deployments, “organic noise”—produced by human users during natural interactions and not pre-programmable—can be collected. Such data is vital for understanding and modeling real user-side behavioral variations. This requires evaluation protocols capable of handling longer-term, more open-ended tasks and quantifying the agent’s performance degradation and recovery capabilities during sustained interactions.
(4) From Phenomenological Description to Theoretical Exploration of the Intrinsic Mechanisms of Robustness The current study primarily documents the phenomenon of performance degradation and some correlative findings (e.g., the orthogonality between reasoning capability and noise robustness). Deeper scientific questions remain: How exactly does noise affect the internal computational processes of an agent? Future research needs to integrate interpretability analysis tools (e.g., neuron activation analysis, attention pattern tracing) to scrutinize the “breaking points” in the agent’s reasoning chain under noise interference. For example, does tool noise primarily disrupt the maintenance of working memory? Does user noise interfere with the parsing of initial conditions for task planning? Understanding these mechanisms is key to designing more targeted interventions. Furthermore, a more solid theoretical framework is needed to formally define agent robustness. This may involve borrowing concepts like stability from control theory or channel capacity against interference from information theory, providing tools for analyzing the upper and lower bounds of robustness for agents with different architectures and training methodologies.
Research on agent noise robustness should not stop at building a more challenging “stress test” benchmark. Its long-term goal is to steer the entire field from pursuing “optimal lab performance” towards constructing “trustworthy agents for the real world.” This demands that we not only evaluate agents in noise but also train agents in noise and understand agents through theory, ultimately achieving a robust symbiosis between agents and the imperfect, uncertain real world. This study is merely the starting point of this journey, and we anticipate more researchers will join this interdisciplinary area, which holds both significant practical value and profound theoretical significance.